# MaxMin-RLHF: Towards Equitable Alignment of Large Language Models with Diverse Human Preferences

### v1

@@ Appendix C Preliminary Results
The parametrized preference probability distribution
$p_{\phi}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})=\frac{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))}{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))+\exp(r_{\phi}(\mathbf{y}_{2},\mathbf{x}))}$
under the Bradley -Terry model (Bradley & Terry, 1952) is Lipschitz with respect to parameter $\phi$. This implies that
$|p_{\phi}-p_{\phi^{\prime}}|\leq L_{p}\|\phi-\phi^{\prime}\|,$ (22)
with $L_{p}=4D$ where $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$.
@@
$\phi^{*}:=\arg\min_{\phi}-\mathbb{E}\big{[}\sum_{\mathbf{z}}p_{\phi^{*}}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})\big{]}$
. As discussed in Sec. 3.2, due to human user groups, a user group specific $\phi^{*}_{u}$ will also exist. Our goal is to characterize the gap between $\hat{\phi}_{\texttt{MLE}}$ and $\phi^{*}_{u}$ defined as
@@
$\phi_{u}^{*}:=\arg\min_{\phi}-\mathbb{E}\big{[}\sum_{\mathbf{z}}p_{u}^{*}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})\big{]}.$ (36)
@@
TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $\leq 2D\|\phi_{u}^{*}-\phi^{*}\|.$ (39)
From the lower bound in (38) and the expression in (37), we obtain
$\|\Delta_{u}^{r}\|\geq\frac{1}{2D}\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}}).$ (40)
Next, we expand on the term TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$. From the statement of Lemma C, we note that $p_{\phi^{*}}({\mathbf{z}})=\sum_{u=1}^{|\mathcal{U}|}\eta(u)p_{\phi_{u}^{*}}({\mathbf{z}})$, hence we can write
TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $=\sum_{\mathbf{z}}(p_{\phi_{u}^{*}}({\mathbf{z}})-p_{\phi^{*}}({\mathbf{z}}))$
$=\sum_{\mathbf{z}}\Big{(}p_{\phi_{u}^{*}}({\mathbf{z}})-\sum_{k=1}^{|\mathcal{U}|}\eta(k)p_{\phi_{k}^{*}}({\mathbf{z}})\Big{)},$ (41)
where we use $k$ to denote a user group for ease of notation. Since $\sum_{k=1}^{|\mathcal{U}|}\eta(k)=1$, we can write
TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $=\sum_{\mathbf{z}}\Big{(}\sum_{k=1}^{|\mathcal{U}|}\eta(k)p_{\phi_{u}^{*}}({\mathbf{z}})-\sum_{k=1}^{|\mathcal{U}|}\eta(k)p_{\phi_{k}^{*}}({\mathbf{z}})\Big{)}.$ (42)
After interchanging the order of summation, we get
TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $=\sum_{k=1}^{|\mathcal{U}|}\eta(k)\Big{(}\sum_{\mathbf{z}}(p_{\phi_{u}^{*}}({\mathbf{z}})-p_{\phi_{k}^{*}}({\mathbf{z}}))\Big{)}$
$=\sum_{k=1}^{|\mathcal{U}|}\eta(k)\cdot\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}_{k}}).$ (43)
@@
Using the equality in (D) into the right hand side of (40), we obtain
$\|\phi^{*}-\phi_{u}^{*}\|\geq\frac{1}{2D}\sum_{k=1}^{|\mathcal{U}|}\eta(k)\cdot\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}_{k}}).$ (44)
From the Definition 2, we will get the final result. Hence proved.
@@
We can define the alignment gap of RLHF model $\pi_{\text{RLHF}}$ to a specific user group $u$ as
@@
We note that in this specific RLHF setting under the KL-based regularization, the objective $-F_{{\phi}}(\pi)$ satisfies strong convexity w.r.t $\pi$ with strong convexity parameter $\mu=1$, hence it holds that
@@
$=\frac{1}{2L_{\pi}\beta^{2}}\|\Psi,\langle\phi^{*}-\phi^{*}_{u},\rangle\|^{2}.$
@@
Hence proved.

### v2

@@ Appendix C Preliminary Results
The parametrized preference probability distribution $p_{\phi}(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})=\frac{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))}{\exp(r_{\phi}(\mathbf{y}_{1},\mathbf{x}))+\exp(r_{\phi}(\mathbf{y}_{2},\mathbf{x}))}$ under the Bradley -Terry model (Bradley & Terry, 1952) is Lipschitz with respect to parameter $\phi$. This implies that
$|p_{\phi}(\mathbf{z})-p_{\phi^{\prime}}(\mathbf{z})|\leq L_{p}\|\phi-\phi^{\prime}\|,$ (15)
where $\mathbf{z}:=(\mathbf{y}_{1}\succ\mathbf{y}_{2}\mid\mathbf{x})$, $L_{p}=4D$, and $D$ denotes the upper bound on the feature representation $\|\psi({\mathbf{y}},{\mathbf{x}})\|\leq D$ for all $({\mathbf{x}},{\mathbf{y}})$.
@@
$\phi^{*}:=\arg\min_{\phi}-\mathbb{E}\bigg{[}\sum_{\mathbf{z}}p_{\phi^{*}}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})\bigg{]}.$
As discussed in Sec. 3.2, due to the presence of diverse human user groups, a $\phi^{*}_{u}$ which, is user group specific, will also exist. Our goal is to characterize the gap between $\hat{\phi}_{\texttt{MLE}}$ and $\phi^{*}_{u}$ defined as
@@
$\phi_{u}^{*}:=\arg\min_{\phi}-\mathbb{E}\bigg{[}\sum_{\mathbf{z}}p_{u}^{*}({\mathbf{z}})\log p_{\phi}({\mathbf{z}})\bigg{]}.$ (29)
@@
TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $\leq 4D\|\phi_{u}^{*}-\phi^{*}\|,$ (32)
where the multiplication of $2$ comes from the fact that there are two terms in the summation in the right side of (31) (cf. Sec. 3.2). From the lower bound in (31) and the expression in (30), we obtain
$\|\Delta_{u}^{r}\|\geq\frac{1}{4D}\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}}).$ (33)
Next, to obtain a lower bound on the term TV ($\text{TV }(p_{\phi_{u}^{*}},p_{\phi^{*}})$, we begin with the definition of the total variation distance $TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})$ as
$TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})$ $=\frac{1}{2}\sum_{z}|p_{\phi_{u}^{*}}(z)-p_{\phi_{j}^{*}}(z)|$
$=\frac{1}{2}\sum_{z}|p_{\phi_{u}^{*}}(z)-p_{\phi^{*}}(z)+p_{\phi^{*}}(z)-p_{\phi_{j}^{*}}(z)|$
$\leq\frac{1}{2}\sum_{z}|p_{\phi_{u}^{*}}(z)-p_{\phi^{*}}(z)|+\frac{1}{2}\sum_{z}|p_{\phi^{*}}(z)-p_{\phi_{j}^{*}}(z)|$
$=TV(p_{\phi_{u}^{*}},p_{\phi^{*}})+TV(p_{\phi_{j}^{*}},p_{\phi^{*}}).$ (34)
In (D), the first equality holds from the definition of TV norm distance, the second equality holds because we add subtract $p_{\phi^{*}}(z)$ inside the norm, we used triangle inequality for the third inequality, and then again utilize the definition of TV distance to write the final equality. After rearranging the terms in (D), we can write
$TV(p_{\phi_{u}^{*}},p_{\phi^{*}})\geq TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-TV(p_{\phi_{j}^{*}},p_{\phi^{*}}).$ (35)
From the definition of $p_{\phi^{*}}(z)=\sum_{k=1}^{|\mathcal{U}|}n(k)p_{\phi_{k}^{*}}(z)$ and from the property of TV distance and Jensen’s inequality, we know that $TV(p_{\phi_{j}^{*}},p_{\phi^{*}})\leq\sum_{k=1}^{|\mathcal{U}|}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}})$. This would imply that $-TV(p_{\phi_{j}^{*}},p_{\phi^{*}})\geq-\sum_{k=1}^{|\mathcal{U}|}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}})$, which we utilize in the right hand side of (35) to obtain
$TV(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $\geq TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-\sum_{k=1}^{|\mathcal{U}|}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}})$
$=(1-\eta(u))TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-\sum_{k\neq u}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}}),$ (36)
which demonstrates the final expression of our lower bound on the term $TV(p_{\phi_{u}^{*}},p_{\phi^{*}})$, which holds in general. Considering the second term in the right-hand side of (D), we can write
$\sum_{k\neq u}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}})\leq TV(p_{\phi^{*}_{k_{\max}(j)}},p_{\phi_{j}^{*}})\sum_{k\neq u}\eta(k),$ (37)
where $TV(p_{\phi^{*}_{k_{\max}(j)}},p_{\phi_{j}^{*}}):=\max_{k\neq u}TV(p_{\phi^{*}_{k}},p_{\phi_{j}^{*}})$ for a given $j$. From the definition of weights $\eta$, we know that $\sum_{k\neq u}\eta(k)=(1-\eta(u))$, hence we can write
$\sum_{k\neq u}\eta(k)TV(p_{\phi_{k}^{*}},p_{\phi_{j}^{*}})\leq(1-\eta(u))TV(p_{\phi^{*}_{k_{\max}(j)}},p_{\phi_{j}^{*}}).$ (38)
@@
Utilizing the upper bound of (38) into the right hand side of (D), we obtain
$TV(p_{\phi_{u}^{*}},p_{\phi^{*}})$ $\geq(1-\eta(u))\Big{[}TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-TV(p_{\phi^{*}_{k_{\max}(j)}},p_{\phi_{j}^{*}})\Big{]}.$ (39)
In the above expression, we note from the right-hand side that the lower bound in (39) holds for all $u$ and $j$. This implies that the right-hand side in (39) will be either positive or negative for $(u,j)$ pairs. But an interesting point to note here is that if the right-hand side is lower bounded away from zero, even for one $(u,j)$ pair, that is an impossibility result, for the corresponding $u$ which is the minority user. To proceed, as a first step, we find the most diverse user-group ($i,j\in\mathcal{U}$) defined as $(i^{*},j^{*})$ be the pair of users with the maximum total variation distance, given by
$(i^{*},j^{*})=\arg\max_{i,j\in\mathcal{U}}TV(p_{\phi_{i}},p_{\phi_{j}}).$ (40)
We next define the minority user group $u^{*}$ as the group with the largest total variation distance from the remaining user groups $k\in\mathcal{U},k\neq j^{*}$, where we quantify this by computing the maximum distance to any other user-groups in the set $\mathcal{U}$. Assuming there exists an
$u^{*}=\arg\max\bigg{\{}\max_{k\neq j}\text{TV}(p^{*}_{i^{*}},p^{*}_{k}),\max_{k\neq i}\text{TV}(p^{*}_{j^{*}},p^{*}_{k})\bigg{\}}.$ (41)
Thus in this process, we select that user-group which is at maximum distance from the set $\mathcal{U}$ and denote as the minority user-group $u^{*}$. Assuming uniqueness in the minority group definition i.e we assume that there exists an unique user-group which is at a maximum distance to the group $j$, we have that $TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})>TV(p_{\phi^{*}_{k_{\max}(j)}},p_{\phi_{j}^{*}})$ for $j=j^{*}$. Using the equality on the right hand side of (33), we obtain
$\|\phi^{*}-\phi_{u}^{*}\|\geq\frac{\epsilon(1-\eta(u))}{4D},$ (42)
where $\epsilon:=TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-\max_{k\neq u}TV(p_{\phi^{*}_{k}},p_{\phi_{j}^{*}})>0$ (for $j=j^{*}$). Hence proved.
@@
We can define the alignment gap of RLHF model $\pi^{*}_{\text{RLHF}}$ to a specific user group $u$ as
@@
We note that in this specific RLHF setting under the KL-based regularization, the objective $-F_{r_{\phi}}(\pi)$ satisfies strong convexity w.r.t $\pi$ with strong convexity parameter $\mu=1$, hence it holds that
@@
$=\frac{1}{2L_{\pi}\beta^{2}}\|\langle\Psi,\phi^{*}-\phi^{*}_{u}\rangle\|^{2}.$
@@
where $\epsilon:=TV(p_{\phi_{u}^{*}},p_{\phi_{j}^{*}})-\max_{k\neq u}TV(p_{\phi^{*}_{k}},p_{\phi_{j}^{*}})>0$. Hence proved.
@@ Appendix F Additional Details of the Experiments
Appendix G Additional Experiments in Robotics Navigation Tasks
In this section, we show that the proposed ideas are well extendable to reinforcement learning in general. We show the performance of MaxMin alignment with single reward RLHF on simple gridworld navigation in Figure 7.
Figure 7: (a) This figure shows a GridWorld navigation scenario where say a government supported vehicle which needs to distribute goods among the two groups denoted by green and orange boxes. In the above figure, (b) shows the trajectory when only green user preferences are considered to decide the vehicle path. (c) shows the trajectory when only green user preferences are considered to decide the vehicle path. (d) shows the result for our proposed formulation, where our goal is to maximize the social utility, which makes sure to develop a robust solution to satisfy all user preferences.